# Hyper-Compression: Model Compression via Hyperfunction

### v1

@@
Feng-Lei Fan E-mail: zeng@math.cuhk.edu.hk Juntong Fan E-mail: zeng@math.cuhk.edu.hk Dayang Wang E-mail: zeng@math.cuhk.edu.hk Jingbo Zhang E-mail: zeng@math.cuhk.edu.hk Zelin Dong E-mail: zeng@math.cuhk.edu.hk Shijun Zhang E-mail: zeng@math.cuhk.edu.hk Ge Wang E-mail: zeng@math.cuhk.edu.hk Tieyong Zeng E-mail: zeng@math.cuhk.edu.hk
(September 1, 2024)
@@
The rapid growth of large models’ size has far outpaced that of GPU memory. To bridge this gap, inspired by the succinct relationship between genotype and phenotype, we turn the model compression problem into the issue of parameter representation to propose the so-called hyper-compression. The hyper-compression uses a hyperfunction to represent the parameters of the target network, and notably, here the hyperfunction is designed per ergodic theory that relates to a problem: if a low-dimensional dynamic system can fill the high-dimensional space eventually. Empirically, the proposed hyper-compression enjoys the following merits: 1) Preferable compression ratio; 2) No post-hoc retraining; 3) Affordable inference time; and 4) Short compression time. It compresses LLaMA2-7B in an hour and achieves close-to-int4-quantization performance, without retraining and with a performance drop of less than 1%. Our work has the potential to invigorate the field of model compression, towards a harmony between the scaling law and the stagnation of hardware upgradation.
@@ 1 Introduction
Recently, the escalating demand for memory and computational resources by large models has presented formidable challenges for their deployment in resource-constrained environments [7], particularly with the stagnation of hardware Moore’s Law [22]. Besides technical obstacles, the substantial cost associated with deployment also engenders profound ethical concerns. The dominance of major institutions in the realm of large models introduces risks pertaining to integrity, bias, security, and other critical aspects. Thus, how to serve large models well is not only a technological imperative but also a pursuit for social good.
@@
One prevalent way to serve large models is to develop effective model compression approaches that crop the size of large models while maintaining acceptable performance levels. Currently, the landscape of model compression predominantly revolves around four fundamental algorithms: pruning, quantization, knowledge distillation, and low-rank approximation [29] whose basic frameworks have been established years ago. However, the compression efficacy of these methods is either capped or challenging to scale. For instance, even the most extreme case of quantization, represented by 1-bit integers, remains upset because the compression ratio is at best a constant.
@@
Here, we introduce a novel and general-purpose approach that redefines model compression as a problem of parameter representation. This concept stems from the principles of hyper-networks, wherein smaller networks can generate weights for a significantly larger target network. We extend the concept of hypernets into what we term a ‘hyperfunction’ (HF), that a small parametric function can generate weights of a large network to advocate hyper-compression. Mathematically, we use a parametric function $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network, and $\Theta$ is the hyperparameters of $h$. Should the memory footprint of $\Theta$ be significantly less than that of $W$, the function $h$ can be stored on devices, allowing for the dynamic querying of weights during inference and thus facilitating the efficient deployment of large-scale models.
@@
One problem that ergodic theory investigates is if a low-dimensional dynamic system can eventually visit a high-dimensional space given sufficient time. We empirically find that the ergodic theory indicates a suitable hyperfunction and leads to a performant hyper-compression algorithm that offers several distinct advantages, succinctly summarized as PNAS: 1) Preferable compression ratio; 2) No post-hoc retraining; 3) Affordable inference time; and 4) Short compression time. More favorably, our method is compatible with other compression methods such as pruning to further escalate the compression efficacy. Lastly, we underscore that model compression via ergodic theory is a novel view whose intrinsic mechanism relates to chaotic theory, Kolmogorov theorem, Diophantine representation problem, etc., which remains unclear. Exploring the underlying compression mechanism will be our future work.
2 Result and Analysis
@@
This section illustrates the efficacy of our novel compression methodology on three widely utilized large, middle, and small models: LlaMA2-7B [21], UNet [18], and MobileNetV3 [13]. Distinctively, our compression technique does not necessitate post-hoc retraining, even at high compression ratios and small performance drop. For example, we achieve up to a 7.93×7.93$\times$7.93 × reduction in the model size of UNet within $1\%$ performance drop. This is a significant improvement over traditional methods that often require extensive retraining to restore efficacy. This capability to compress without much loss of performance potentially sets a new benchmark in model compression. Detailed descriptions of our algorithm and experiments are provided in Supplementary Information.
@@
Table 1: The compression effectiveness of our method on UNet, MobileNetV3 and LlaMA2-7B. As a reference, the compression ratio of INT4 quantization is 4×4$\times$4 ×.
@@
HF 6.45 (7.93$\times$) 99.71
Pruning + HF 2.86 (17.87$\times$) 96.45
@@
HF 1.18 (5.04$\times$) 73.93
Pruning + HF 0.43 (13.84$\times$) 68.47
Table 2: Changes in inference time on UNet
Batch Size 1 2 4 8 16
Normal-HF (s) 5.70 5.88 6.00 7.36 15.64
Optimized-HF (s) 1.17 1.22 1.39 2.70 11.12
Original (s) 0.18 0.28 0.48 1.78 10.07
@@
Preferable compression ratio. As shown in Table 4, we apply the same pruning method used in Sheared-LlaMA-1.3B, TinyLlaMA, and LiteLlaMA, combined with our compression technique. We evaluate our compressed models on eight downstream tasks: 0-shot accuracy on SciQ, WinoGrande, ARC-E, 25-shot ARC-C, 10-shot HellaSwag, 32-shot BoolQ, NQ, and 5-shot MMLU. The average test results of eight downstream tasks are shown in Table 4, where “File Size" indicates the size of the model parameter file. The comparison of individual tasks is put in Supplementary Information. It is noteworthy that our method can compress LlaMA2-7B by a factor of 2.60×2.60$\times$2.60 × while maintaining the Average score decrease within 1%, whereas Sheared-LlaMA-1.3B achieves 2.60×2.60$\times$2.60 × with $15.57\%$ performance decreases.
As for UNet and MobileNetV3, our method can compress UNet and Pruned-UNet by 7.93×7.93$\times$7.93 × and 7.10×7.10$\times$7.10 × with the performance loss contained in 1%. Particularly, our method succeeds in combination with other model compression methods such as pruning to achieve an even higher compression ratio. In UNet, the total compression ratio is 17.87×17.87$\times$17.87 × with the performance loss $3.41\%$.
@@
No post-hoc retraining. Despite the presence of minor discrepancies between the original and decoded parameters—specifically, at magnitudes ranging from $10^{-4}$ to $10^{-3}$—the impact on error accumulation through layer-by-layer information propagation remains acceptable. This is due to the inherent redundancy within the network itself, including parametric redundancy and activation space redundancy. As a result, the destructive effect of our method is minimal, obviating the need for post-hoc retraining. This characteristic is particularly advantageous in real-world settings, where often 1) training data are inaccessible, and curating data for fine-tuning is costly; 2) no computing resources are supplied for retraining.
@@
Affordable inference time. A network is a hierarchical structure. To expedite the inference time, our approach is to parallelize the decoding and inference processes. The key is to complete the decoding of a layer before using this layer to infer. As shown in Table 2, “Normal-HF" means the model without decoding and inference parallized. "Original" refers to the time required for a single inference of the original UNet, and "Optimized" refers to the time required for one inference by employing parallel decoding and inference processes. when batch size increases, the inference time of our compressed model gets closer to that of the original model, which aligns with our expectations. This is because when batch size goes up, the inference time across layers is slow, which allows weights in later layers to be decompressed in time.
@@
Short compression time. As Table 6 shows, our method can compress models very fast. This efficiency primarily stems from matrix operations to expedite the majority of the compression process and the K-D tree to expedite the search. Additionally, by treating the compression tasks between layers as independent operations, we implement a multiprocessing parallel strategy to further decrease compression times.
@@
3 Materials and Methods
Motivation
A human genome of only approximately 30,000 genes can remarkably encode the development of a human brain that has finally billions of neurons and trillions of synaptic connections [20]. This observation suggests the existence of a concise and implicit mapping from genotype to phenotype, capable of expanding a limited number of genetic instructions into a vast array of biological functionalities. Inspired by this efficient genetic encoding, the idea of hypernet is to use a small network (genotype, called hypernet) to generate weights for the target network (phenotype).
We consider the hypernet in the setting of model compression. Furthermore, we generalize the idea of hypernet to a parameterized function (hyperfunction). The biological observation is the existence of a concise mapping, regardless of whether it is denoted by a network or other forms of functions. With a hyperfunction, instead of directly diminishing weights, one stores a hyperfunction than the model and queries its weights when needed. This novel perspective converts the model compression problem into the parameter representation issue. Mathematically, we use a hyperfunction $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network, and $\Theta$ is the hyperparameters of $h$. Along this direction, the important question is how can we design a suitable hyperfunction?
@@ 3 Materials and Methods
The above question can be affirmatively answered through the lens of ergodic theory. One branch of the ergodic theory [6] investigates if the trajectory of a low-dimensional dynamical system will eventually traverse every point of a given high-dimensional space. From the engineering point of view, the ergodic theory discusses the feasibility of using low-dimensional curves to approximate points in high-dimensional space, which indicates a potentially ideal hyperfunction for model compression.
@@
For any $K\in\mathbb{N}^{+}$, suppose $\alpha_{1},\cdots,\alpha_{K}$ are irrationally independent, the following point set
$\left\{\left[\tau\left(a_{1}\cdot\theta\right),\tau\left(a_{2}\cdot\theta\right),\cdots,\tau\left(a_{K}\cdot\theta\right)\right]^{T}:\theta\in[0,+\infty]\right\}$ (1)
is dense in $[0,1]^{K}$, where $\tau(x)=x-[x]$.
In Theorem 1, the irrationality condition is easily fulfilled. For example, we can set $a_{k}=1/(\pi+k)$. Given $\{W_{n}\}_{n=1}^{N}$, according to Theorem 1, we know that for any $\epsilon>0$, there exists $\theta^{*}$ such that
$|W_{n}-\tau(\theta^{*}/(\pi+n))|<\epsilon,n=1,\cdots,N.$ (2)
Theorem 1 provides a natural and elegant way to encode a group of numbers into one number. However, we should only be cautiously optimistic about it since simply stitching digits of numbers can also ensemble a group of numbers into one number. What matters is whether we can find a simple $\theta$ that approximates $W_{1},\cdots,W_{K}$ based on (2), such that the memory footprint of the former is smaller than the latter. Our experiments empirically confirm that the parametric representation according to the ergodic theory has compression efficacy. This is primarily because our compression is lossy, and the network itself has redundancy.
@@
The entire algorithmic development is anchored in (2). It begins by partitioning the parameters $\{W_{n}\}_{n=1}^{N}$ into groups, each with $K$ parameters. The core objective of the algorithm is, given $\alpha_{1},\cdots,\alpha_{K}$ and an acceptable error $\epsilon$, to identify the smallest $\theta^{*}$ for points in $[0,1]^{K}$, aiming for the highest feasible dimension $K$. First, though the existence of $\theta^{*}$, because a network usually has millions or billions of parameters, searching $\theta^{*}$ for a group of $K$ parameters has to be fast. Second, an imprecise $\theta^{*}$ could lead to significant performance degradation in the restored networks. Therefore, the algorithm must ensure that $\theta^{*}$ approximates the parameters precisely. Third, when time and precision allow, the algorithm should return the smallest $\theta^{*}$ for the highest compression ratio. The algorithm details are shown in Supplementary Information.
4 Conclusion and Future Work
In this study, we have proposed hyper-compression, a novel and general-purpose methodology for model compression that leverages hyperfunction to encode the parameters of the target network. We have conducted experiments on various architectures, including LlaMA2-7B, UNet, and MobileNetV3, demonstrating that our approach is both user-friendly and scalable. Nevertheless, it represents merely an initial foray into the whole blueprint. Future directions include 1) theoretically describing the relationship among the error, the length of $\theta$, and the dimension $K$ to reveal the essence of using the ergodic theory to realize compression; 2) beyond classical theory, exploring other possibilities such as implicit neural representation [3] to strike a better balance between compression time, inference time, and compression ratio. We believe that hyper-compression can contribute to Moore’s law of model compression in the near future, i.e., the compression efficiency of model compression algorithms can be doubled annually, to meet the need of large models’s development.
@@ 5 Supplementary Information
I. Limit of the Compression Ratio of Pruning Convolutional Networks
In the main body, we mention that pruning-based methods are hard to scale. Here, we show that the compression ratio of pruning a convolutional network is constantly bounded by studying a randomly generated infinite network and introducing the percolation theory [10] that characterizes the connectivity of a network when links are removed per certain rules.
We study one-dimensional convolutional networks whose topology is shown in Figure 5. We assume any edge of a network is randomly removed with the same probability. This is also equivalent to the following process: Randomly generate weights for edges of a given network. If the weight of an edge is higher than a threshold, the edge is retained; otherwise, it is removed. In the jargon of percolation theory, an edge is open or closed if the edge exists or disappears. When the probability of removing edges reaches a threshold, referred to as the percolation threshold, there will be no connected clusters from the input and output, meaning that the network will always yield 00 for any input. The probability threshold indicates the maximal proportion of a network’s edges to prune. This section discusses the upper and lower bounds of the percolation threshold for a class of graphs defined by the one-dimensional convolutional network, thereby quantitatively characterizing the compression ratio.

### v2

@@
Feng-Lei Fan E-mail: zeng@math.cuhk.edu.hk Juntong Fan E-mail: zeng@math.cuhk.edu.hk Dayang Wang Feng-Lei Fan, Juntong Fan, and Dayang Wang contribute equally. E-mail: zeng@math.cuhk.edu.hk Jingbo Zhang E-mail: zeng@math.cuhk.edu.hk Zelin Dong E-mail: zeng@math.cuhk.edu.hk Shijun Zhang E-mail: zeng@math.cuhk.edu.hk Ge Wang E-mail: zeng@math.cuhk.edu.hk Tieyong Zeng E-mail: zeng@math.cuhk.edu.hk
(December 14, 2024)
@@
The rapid growth of large models’ size has far outpaced that of GPU memory. To bridge this gap, inspired by the parsimonious relationship between genotype and phenotype, we turn the model compression problem into the issue of parameter representation to propose the so-called hyper-compression. The hyper-compression uses a hyperfunction to represent the parameters of the target network per ergodic theory, that addresses the following approximation problem: if a low-dimensional dynamic system can fill the high-dimensional space eventually. Empirically, the proposed hyper-compression enjoys the following merits: 1) Preferable compression ratio; 2) No post-hoc retraining; 3) Affordable inference time; and 4) Short compression time. It compresses LLaMA2-7B in an hour and achieves close-to-int4-quantization performance, without retraining and with a performance drop of less than 1%. Our work can facilitate the harmony between the scaling law and the stagnation of hardware upgradation in terms of saving both computation and data. We have open-sourced our code for readers’ free download and evaluation.
@@ 1 Introduction
Recently, due to the pursuit of the scaling law, the escalating demand for memory and computational resources by large models has presented formidable challenges for their deployment in resource-constrained environments [8], particularly with the stagnation of hardware Moore’s Law [25]. Besides technical obstacles, the substantial cost associated with deployment also causes the dominance of major institutions in the community, which introduces risks of integrity, bias, security, and other critical aspects. Thus, how to serve large models well is not only a technological imperative but also a strategic pursuit for social good.
@@
One prevalent way to serve large models is to develop effective model compression approaches that crop the size of large models while maintaining the performance of the original model. Currently, the landscape of model compression predominantly revolves around four different classes of algorithms: pruning, quantization, knowledge distillation, and low-rank decomposition [34] whose fundamental frameworks were established years ago. However, the compression efficacy of these methods is challenging to scale. For instance, even the most extreme case of quantization, represented by 1-bit integers, remains unexciting because the compression ratio is at best a constant, which is handicapped to address the enlarged gap between models and the hardware resource. Moreover, when compressing a model a lot, the model’s output is severely distorted. Thus, retraining and recalibration are mandated to recover the model’s performance.
@@
Here, we introduce hyper-compression, a novel and general-purpose approach that redefines model compression as a problem of parameter representation. This concept stems from hyper-networks that are smaller networks but can generate weights for a significantly larger target network, in analogy to the biological observation that a comparatively small genome can encode a large neuronal network. We generalize the concept of hypernets into what we term a ‘hyperfunction’ (HF) and hypothesize that weights of a large network can be encoded into a small parametric function. Mathematically, we use a parametric function $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network, and $\Theta$ is the hyperparameters of $h$. Should the memory footprint of $\Theta$ be significantly less than that of $W$, the function $h$ can be stored on devices, allowing for the dynamic hierarchical querying of weights during inference and thus facilitating the efficient deployment of large-scale models.
@@
One of the most important problems that ergodic theory investigates is if a low-dimensional dynamic system can eventually visit a high-dimensional space given sufficient time. We empirically find that the ergodic theory indicates a suitable hyperfunction and leads to a performant hyper-compression algorithm that offers several distinct advantages, succinctly summarized as PNAS: 1) Preferable compression ratio; 2) No post-hoc retraining&recalibration; 3) Affordable inference time; and 4) Short compression time. More favorably, our method is compatible with other compression methods such as pruning to further escalate the compression efficacy multiplicatively. Lastly, we underscore that model compression via ergodic theory is a novel view whose intrinsic mechanism, relating to chaotic theory, Kolmogorov theorem, Diophantine representation problem, etc., remains unclear. Given the criticality of model compression, we report our empirical results first, and exploring the mechanism of hyper-compression will be our future work.
Result and Analysis
@@
This section illustrates the efficacy of our novel compression methodology on three widely utilized large, middle, and small models: LlaMA2-7B [24], UNet [20], and MobileNetV3 [15]. Distinctively, our compression technique does not necessitate post-hoc retraining, even at high compression ratios. For example, we achieve up to a 7.87×7.87$\times$7.87 × reduction in the model size of UNet within $1\%$ performance drop. This is a significant improvement over traditional methods that often require extensive retraining to restore efficacy. This capability sets a new benchmark in model compression. Detailed descriptions of our algorithm and experiments are provided in Supplementary Information.
@@
Table 1: The compression effectiveness of our method on UNet, MobileNetV3, and LlaMA2-7B. As a reference, the compression ratios of INT4 quantization in LlaMA-7B, UNet, and MobileNetV3 are 4×4$\times$4 ×, 8×8$\times$8 ×, and 8×8$\times$8 ×, respectively.
@@
HF 6.49 (7.87$\times$) 99.71
Pruning + HF 2.88 (17.74$\times$) 96.45
@@
HF 1.65 (3.61$\times$) 73.93
Pruning + HF 0.47 (12.66$\times$) 68.47
Batch Size 4060 A40
Original (s) Ours (s) Original (s) Ours (s)
1 4.35 4.43 1.50 2.06
2 3.95 4.49 1.44 2.05
4 3.63 4.77 1.45 2.22
8 3.67 7.94 1.44 2.46
@@
Preferable compression ratio. As shown in Table 4, we apply the same pruning method used in Sheared-LlaMA-1.3B, TinyLlaMA, and LiteLlaMA, combined with our compression technique. We evaluate our compressed models on eight downstream tasks: 0-shot accuracy on SciQ, WinoGrande, ARC-E, 25-shot ARC-C, 10-shot HellaSwag, 32-shot BoolQ, NQ, and 5-shot MMLU. The average test results of eight downstream tasks are shown in Table 4, where “File Size" indicates the size of the file that stores the model. The comparison of individual tasks is put in Supplementary Information. Notably, our method can compress LlaMA2-7B by a factor of 2.60×2.60$\times$2.60 × while maintaining the average score decrease within 1%, whereas Sheared-LlaMA-1.3B achieves 2.50×2.50$\times$2.50 × with $15.57\%$ performance decreases. As for UNet and MobileNetV3, our method can compress UNet and Pruned-UNet by 7.87×7.87$\times$7.87 × and 7.05×7.05$\times$7.05 × with the performance loss contained in 1%. Particularly, our method succeeds in combination with other model compression methods such as pruning to achieve an even higher compression ratio. In UNet, the total compression ratio is 17.74×17.74$\times$17.74 × with the performance loss $3.41\%$.
@@
No post-hoc retraining. Because discrepancies between the original and decoded parameters are minor, at magnitudes ranging from $10^{-4}$ to $10^{-3}$, the impact on error accumulation through layer-by-layer propagation is acceptable. As a result, the destructive effect of our method is minimal, obviating the need for post-hoc retraining. This advantage is particularly useful in real-world settings, where often 1) training data are inaccessible, and curating data for fine-tuning is costly; 2) no computing resources are supplied for retraining.
@@
Affordable inference time. A network is a hierarchical structure. To expedite the inference time, our approach is to leverage this hierarchy by parallelizing the decoding and inference processes. The key is to complete the decoding of a layer before using this layer to infer. As shown in Table 2, “Original" refers to the time required for a single inference of the original UNet, and “Ours" refers to the time required for one inference by employing parallel decoding and inference processes. It can be seen that the hyper-compression only increases the total inference time moderately.
@@
Short compression time. As Table 6 shows, our method can compress models very fast. This efficiency primarily stems from using matrix operations for most compression processes and the K-D tree to expedite the search. Additionally, by treating the compression tasks between layers as independent operations, we implement a multiprocessing parallel strategy to further decrease the compression time.
Materials and Methods
Motivation
A small human genome can remarkably encode the development of a human brain that has finally billions of neurons and trillions of synaptic connections [23, 22]. This observation suggests the existence of an implicit mapping from genotype to phenotype, capable of expanding a limited number of genetic instructions into a vast array of biological substances. Inspired by this efficient genetic encoding, the hypernet uses a small network (genotype, called hypernet) to generate weights for the target network (phenotype).
@@
We consider the hypernet in the setting of model compression. Furthermore, we generalize the idea of hypernet to a parameterized function (hyperfunction). The biological observation is the existence of a mapping, regardless of whether it is denoted by a network or other forms of functions. With a hyperfunction, instead of directly diminishing weights, one stores a hyperfunction about the model and queries its weights when needed. This novel perspective converts the model compression problem into the parameter representation issue. Mathematically, we use a hyperfunction $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network, and $\Theta$ is the hyperparameters of $h$. Along this direction, the important question is how can we design a suitable hyperfunction?
@@
The above question fits the scope of Diophantine approximations or ergodic theory. A branch of ergodic theory [6] studies whether a low-dimensional system’s trajectory can cover all points in a high-dimensional space. From an engineering perspective, it assesses the feasibility of using low-dimensional curves to approximate high-dimensional points, indicating the potential for model compression. The following theorem illustrates a simple yet notable example from ergodic theory.
@@
Suppose $a_{1},\cdots,a_{N}$ are irrationally independent, for any given set $\{W_{n}\}_{n=1}^{N}\subseteq[0,1]$ and $\epsilon>0$, there exists a value $\theta^{*}\in[0,+\infty)$ such that
$|W_{n}-\tau(\theta^{*}a_{n})|<\epsilon,\;n=1,\cdots,N,$ (1)
where $\tau(x)=x-\lfloor x\rfloor$.
The proof of Theorem 1 is a direct consequence of Lemma 19 from [32]. In this theorem, the irrationality independence condition is readily fulfilled; for instance, one can select $a_{k}=1/(\pi+k)$. Theorem 1 provides a natural and elegant way to encode a group of numbers into one number. However, we should only be cautiously optimistic about its application in compression, since simply stitching digits of numbers can also ensemble a group of numbers into one number. What matters is whether we can find a simple $\theta$ that approximates $W_{1},\cdots,W_{K}$ based on (1), such that the memory footprint of the former is smaller than the latter. Our experiments empirically confirm that the parametric representation according to Theorem 1 has compression efficacy.
@@
The entire algorithmic development is anchored in (1). The core objective of the algorithm is, given $\alpha_{1},\cdots,\alpha_{K}$ and an acceptable error $\epsilon$, to identify the smallest $\theta^{*}$ for points in $[0,1]^{K}$ and the highest feasible dimension $K$. First, because a large model usually has billions of parameters, searching $\theta^{*}$ for a group of $K$ parameters has to be fast. Second, an imprecise $\theta^{*}$ could lead to significant performance degradation in the restored networks. Therefore, the algorithm must ensure that $\theta^{*}$ assists in approximating a network’s parameters precisely. Third, when time and precision allow, the algorithm should return the smallest $\theta^{*}$ for the highest compression ratio.
Conclusion and Future Work
In this study, we have proposed hyper-compression, a novel and general-purpose methodology for model compression that leverages hyperfunction to encode the parameters of the target network. We have conducted experiments on various architectures, including LlaMA2-7B, UNet, and MobileNetV3, demonstrating that our approach is both user-friendly and scalable. Nevertheless, it represents merely an initial foray into the whole blueprint. Future directions include 1) theoretically describing the relationship among the error, the length of $\theta$, and the dimension $K$ to reveal the essence of using the ergodic theory to realize compression; 2) beyond classical theory, exploring other possibilities such as implicit neural representation [3] to strike a better balance between compression time, inference time, and compression ratio. We believe that hyper-compression can contribute to Moore’s law of model compression in the near future, i.e., the compression efficiency can be doubled annually, as a solution for the stagnation of hardware Moore’s law.
@@ 2 Supplementary Information
I. The Enlarged Gap Between Models’ Size and Hardware Capacity
In recent years, the field of AI has witnessed remarkable advancements, particularly with the development of large language models (LLMs). These large models have demonstrated impressive capabilities in various tasks, which was widely believed to make a big stride towards artificial general intelligence. Large models have rapidly increased in size over the past few years. For instance, GPT-3 consists of 175 billion parameters. Recently, PanGu-$\Sigma$, one of the largest models to date, staggeringly has over $10^{3}$ billion parameters. Such a growth in model size has been driven by the desire to improve the model’s complex language understanding by scaling law. On the other hand, the development of GPU computing capacity such as memory has not kept pace. One main reason is that the fabrication of more advanced GPUs gradually gets closer to the physical ceiling. The disparity between models’ size and GPU memory has created significant challenges in the development of large models.
Specifically, the disparity between the model size and GPU capacity brings the following problems:
  1. 1.
    Memory Bottleneck: The limited GPU memory poses a bottleneck for serving LLMs. When deploying LLMs for real-world applications, the entire model and its associated parameters need to fit within the GPU memory. However, as LLMs continue to grow in size, it becomes increasingly difficult to accommodate them within the limited memory constraints. This limitation hampers the deployment and wide adoption of LLMs in the edge settings.
  2. 2.
    Ethic Issues: Since large models are hard to serve, large institutes will dominate the development and deployment of large language models, which upsets the whole AI ecosystem:
    Algorithmic Bias: LLMs developed by dominant institutes may be biased by the data they were trained on. This can perpetuate existing societal biases and discrimination, leading to unfair outcomes in decision-making processes.
    Lack of Transparency: Large institutes may not always disclose the full details of their model architectures, training data, or decision-making processes. The lack of transparency can hinder accountability, making it difficult to assess the fairness and reliability of LLMs in real-world applications.
As the demand for more powerful models and the risk of monopoly by large institutes continue to rise, addressing the disparity between the model size and GPU capacity becomes increasingly important. Particularly, we need to cultivate novel model compression techniques that can realize "Moore’s law of model compression", as a solution for the stagnation of hardware Moore’s law. Once accomplished, Moore’s law of model compression, which proposes that the compression ratios of models should double annually, can constitute a solid infrastructure to support the whole AI ecosystem.
II. Limit of the Compression Ratio of Pruning Convolutional Networks
In the main body, we mention that pruning-based methods are hard to scale. Here, we show that under certain conditions, the compression ratio of pruning a convolutional network is constantly bounded. The network is a randomly generated network with infinite width and depth. We introduce the percolation theory [12] that characterizes the connectivity of a network when links are removed per certain rules.
We study one-dimensional convolutional networks whose topology is shown in Figure 6. We assume any edge of a network is randomly pruned with the same probability. This is also equivalent to the following pruning process: Randomly generate weights for edges of a given network. If the weight of an edge is higher than a threshold, the edge is retained; otherwise, it is removed. In the jargon of percolation theory, an edge is called open or closed if the edge exists or disappears. The percolation theory is a well-known phase transition theory. When the probability of removing edges reaches a threshold, referred to as the percolation threshold, there will be no connected clusters from the input and output, meaning that the network will always yield 00 for any input. The probability threshold directly indicates the maximal proportion of a network’s edges to prune. This section discusses the upper and lower bounds of the percolation threshold for a class of graphs defined by the one-dimensional convolutional network, thereby quantitatively characterizing the compression ratio.